1. One liter of a saturated sol’n of Ag2CrO4 at 25°C contains 0.0435 g of
Ag2CrO4. Calculate its Ksp.

2. Calculate a) molar solubility, b) [ ] of the constituent ions, and c)
solubility in g/l for Zn(OH)2. Ksp = 4.5 x 10-17.
Answer #1

1)

Molar mass of Ag2CrO4= 331.8 g/mol

s = 0.0435 g/L = 4.35*10^-2 g/L

To covert it to mol/L, divide it by molar mass

s = 4.35*10^-2 g/L / 331.8 g/mol

s = 1.311*10^-4 g/mol

The salt dissolves as:

Ag2CrO4 <----> 2 Ag+ + CrO42-

   2s s

Ksp = [Ag+]^2[CrO42-]

Ksp = (2s)^2*(s)

Ksp = 4(s)^3

Ksp = 4(1.311*10^-4)^3

Ksp = 9.01*10^-12

Answer: 9.01*10^-12

Feel free to comment below if you have any doubts or if this answer do not work

Only 1 question at a time please